in a certain lottery, 4 different numbers between 1 and 11 inclusive are drawn. These are the winning numbers. A player selects 4 different number between 1 and 11. What is the probability that the player has all the winning numbers?
Answer by Fombitz(32388)

Between 1 and 11 there are 11 numbers.
Once a number is chosen it can't be chosen again.
P(1st winning number)=1/11
P(2nd)=2/10=1/5
P(3rd)=3/9=1/3
P(4th)=4/8=1/2
P(4 winners)=(1/11)(1/5)(1/3)(1/2)=1/330